### v1

# CFBench: A Comprehensive Constraints-Following Benchmark for LLMs

Abstract
The adeptness of Large Language Models (LLMs) in comprehending and following natural language instructions is critical for their deployment in sophisticated real-world applications. Existing evaluations mainly focus on fragmented constraints or narrow scenarios, but they overlook the comprehensiveness and authenticity of constraints from the user’s perspective. To bridge this gap, we propose CFBench, a large-scale Comprehensive Constraints Following Benchmark for LLMs, featuring 1,000 curated samples that cover more than 200 real-life scenarios and over 50 NLP tasks. CFBench meticulously compiles constraints from real-world instructions and constructs an innovative systematic framework for constraint types, which includes 10 primary categories and over 25 subcategories, and ensures each constraint is seamlessly integrated within the instructions. To make certain that the evaluation of LLM outputs aligns with user perceptions, we propose an advanced methodology that integrates multi-dimensional assessment criteria with requirement prioritization, covering various perspectives of constraints, instructions, and requirement fulfillment. Evaluating current leading LLMs on CFBench reveals substantial room for improvement in constraints following, and we further investigate influencing factors and enhancement strategies. The data and code are publicly available at https://github.com/PKU-Baichuan-MLSystemLab/CFBench.
Figure 1: Sample data from CFBench. A checklist, constraint type, requirement priority, and satisfaction constitute our evaluation criteria.
Introduction
Large Language Models (LLMs) have become the cornerstone of numerous cutting-edge research tasks and are widely utilized in real-world scenarios (Brown et al. 2020; Chowdhery et al. 2023; Achiam et al. 2023; Touvron et al. 2023). In real-world scenarios, human instructions are inherently complex and accompanied by explicit constraints, requiring models to both understand intricate requirements and strictly comply with these constraints (Yang et al. 2023; Zhong et al. 2021; Mishra et al. 2022; Wei et al. 2021; Sanh et al. 2022). The proficiency of LLMs in comprehending requirements and adhering to natural language constraints is essential, as it ensures tasks are executed precisely and resolved perfectly according to user instructions.
The prevailing method for evaluating a model’s instruction-following ability involves using quantitative programs, human evaluators, or advanced LLMs to assess response quality across single constraints, complex problems, and finite constraints  (Zhou et al. 2023a; Wang et al. 2023; Li et al. 2023; Zheng et al. 2024; Xu et al. 2023).  Laskar et al. (2024) underscores the importance of evaluating data quality, highlighting the necessity for real and extensive data distribution, along with its applicability to real-world scenarios.  Sun et al. (2024b) also stresses that realistic evaluation metrics reflect model capabilities and guide iteration. Constraints-following evaluation faces analogous challenges, particularly within complex real-world scenarios, where data sources and contexts are diverse, and where evaluation is both subjective and arduous. Fig. 1, which addresses the aforementioned challenges, presents a sample from CFBench illustrating the Trump assassination event with different colors representing various constraint types. The instruction include multiple constraints, and the evaluation method uses a checklist to break down complex requirements into independent checkpoints, annotating constraint types and priorities. LLMs are then used to assess each checkpoint. Consequently, we introduce two more profound challenges in constraints-following assessment.
Q1: How to construct high-quality evaluation data?
Many studies focus on evaluating single constraint (Chen et al. 2022; Tang et al. 2023), lacking comprehensive analysis across diverse constraints.  He et al. (2024b) examines LLM performance on complex real-world instructions but neglect constraint diversity and scenario coverage. Jiang et al. (2023) incrementally incorporate fine-grained constraints to craft multi-level instructions. However, with only 75 instances of mixed type, which risks variability due to limited data, and equating difficulty with constraint quantity oversimplifies the task. Recent work focuses on evaluating constraints combinability (Wen et al. 2024). To ensure data quality, we systematically categorize constraints by mining real-world online data and using classification, synthesis, and expert design, covering 10 primary categories and over 25 subcategories. We also cross-match these constraints with various domains and scenarios, ensuring balanced representation and expert-validated reasonableness.
Q2:How to evaluate accurately and meticulously?
Evaluating LLMs’ adherence to constraints is challenging and typically involves manual, automated, and programmatic assessments using various metrics. Representative work computes outcomes for verifiable instructions using code  (Zhou et al. 2023a; He et al. 2024b).  Jiang et al. (2023) uses scripts and constraint-evolution paths to handle diverse challenging instructions, introducing three metrics tailored to the data’s characteristics. The DRFR method decomposes complex constraints into binary judgments, with GPT evaluating each criterion (Qin et al. 2024). Indeed, previous work has ensured the feasibility and objectivity of evaluations through various methods, but they have overlooked assessments from the user’s multiple perspectives. We deconstruct complex instructions from the user’s perspective into multiple sub-needs, categorizing them by priority and constraint type, with LLMs evaluating each checkpoint. Furthermore, a multi-dimensional evaluation criteria is proposed using three metrics from the perspectives of constraints, instructions, and requirements priority.
We introduce CFBench, a comprehensive benchmark designed to thoroughly evaluate the constraint comprehension and following capabilities of LLMs. CFBench aggregates constraints from real-world scenarios and pioneers a hierarchical system for constraint types, encompassing 10 primary categories and over 25 secondary subcategories organized through taxonomic and statistical methodologies. CFBench features 1,000 meticulously curated samples spanning more than 200 real-life scenarios across 20 domains and over 50 NLP tasks, enhancing the breadth and generality of the evaluation data. Additionally, we have seamlessly integrated original instructions and constraint types within each sample, paying particular attention to nuanced combinations, ensuring each constraint is credibly and coherently embedded. Our advanced evaluation methodology incorporates multi-dimensional assessment criteria, which prioritizing requirements to align LLM outputs with user needs, enhance interpretability, and facilitate iterative development. Finally, extensive experiments and exploratory discussions provide strong support for evaluation and optimization.
Overall, our contributions are mainly four-fold:
  • To the best of our knowledge, we are the pioneers in systematically defining an instruction constraint framework utilizing both taxonomic and statistical methodologies.
  • We introduce CFBench, a meticulously annotated, large-scale, high-quality benchmark that encompasses a broad spectrum of real-world scenarios and NLP tasks.
  • We propose a multi-dimensional evaluation framework to comprehensively assess model capabilities while prioritizing user-centric needs.
  • We exhaustively evaluated prominent LLMs, uncovering significant deficiencies in constraints following and exploring performance factors and optimization strategies.
Figure 2: The construction pipeline of CFBench. Initially, it entails the construction of the constraint system, followed by the assembly of the dataset, and culminating in the proposal of a multi-perspective user view evaluation.
Related Work
Evaluation for LLMs
Numerous studies have concurrently evaluated LLMs from various perspectives. By integrating a multitude of existing datasets, many have assessed the overall capabilities of LLMs  (Bommasani, Liang, and Lee 2023; Zhong et al. 2024; Dubois et al. 2024; Chia et al. 2024; Hendrycks et al. 2021). Some research has delved into specialized capabilities, such as programming  (Chen et al. 2021), reasoning  (Cobbe et al. 2021), and knowledge  (Huang et al. 2024). Unlike previous studies, we evaluate LLMs’ instruction-following abilities with a focus on constraint-based instructions.
Instruction Following
Fine-tuning LLMs with annotated instructional data enhances their ability to follow general language instructions  (Weller et al. 2020; Sanh et al. 2022; Mishra et al. 2022). Studies show that more complex or constrained instructions further improve this ability. For instance, six methods to create intricate instructions from a small set of handwritten seed data are proposed (Xu et al. 2023; Luo et al. 2023), while  Mukherjee et al. (2023) elevate training data complexity by having GPT-4 generate reasoning steps for simple instructions. The latest work  (Sun et al. 2024a; He et al. 2024a; Dong et al. 2024) suggests that increasing the number and variety of constraints can enhance the complexity of instructions, thereby further improving the model’s ability to follow constraint-based instructions.
Evaluation of Constraints Following
Constraints such as word count, position, topics, and content have garnered significant attention in the field of Controlled Text Generation  (Yao et al. 2023; Zhou et al. 2023b; Chen et al. 2022). Zhou et al. (2023a) centers on assessing 25 verifiable instructions. Numerous studies have explored the adherence of LLMs to format constraints, such as complex tabular data  (Tang et al. 2023) and customized scenario formats  (Xia et al. 2024).  Qin et al. (2024) decomposing a single instruction into multiple constraints. He et al. (2024b) gathered constraints from real-world scenarios and developed a sophisticated benchmark using detailed task descriptions and inputs.  Jiang et al. (2023) progressively integrates fine-grained constraints to develop multi-level instructions, thereby enhancing complexity across six distinct types. Concurrent work  (Wen et al. 2024), constructs a novel benchmark by synthesizing and refining data from the aforementioned benchmarks, with an emphasis on the combinatorial types of constraints. However, previous studies suffered from fragmented constraints, limited scenarios, and misaligned evaluation methods with user perspectives.
Split Set Basic Info Constraints Count
Num. Len. Prim. Cons. Type. C1 C2 C3 C4 C5 C6 C7 C8 C9 C10
Easy Set 500 413 1.69 3.59 2.83 613 214 180 170 134 92 82 95 90 79
Hard Set 500 605 1.98 4.89 3.58 772 345 233 241 168 122 115 145 137 81
Full Set 1000 509 1.84 4.24 3.20 1385 559 413 411 302 214 197 240 227 160
Table 1: CFBench Statistic. The abbreviations of ’Num.’, ’Len.’, ’Prim.’, ’Cons.’, ’Type.’ denote the sample number, average instruction length, average primary requirements number, average constraint number, average constraint type number, respectively. The designations ’C1’-’C10’ denote the Primary Constraint types of content, numerical, style, format, linguistic, situation, example, inverse, contradictory, and rule constraint, respectively.
CFBench
As depicted in Fig. 2, the CFBench construction pipeline includes several key components. First, we collect and systematize constraint expressions from real-world scenarios and various NLP tasks. Using this system, we create high-quality evaluation data by combining instructions from these scenarios with advanced LLMs and manual curation. We then introduce innovative multi-perspective evaluation method. Additionally, we conduct a thorough statistical analysis and validate the quality from various angles to highlight reliability and applicability.
Constraints System
Constraints Collection
We amass a diverse corpus of instructions from real-world scenarios and various NLP tasks  (Xia et al. 2024; Li et al. 2024) to ensure a comprehensive system. Initially, we aggregate several million instructions from online logs and NLP tasks, refining these through length filtering and clustering to distill 30,000 high-quality instructions. Utilizing advanced LLM techniques, we extract and expand atomic constraints through evolutionary methods. Using LLMs, we carefully select meaningful atomic constraints, resulting in over 5000 unique constraints. Domain experts first filter out unreasonable or meaningless atomic constraints and then synthesize these into a structured framework with 10 primary categories and 25 subcategories, guided by principles of statistics, taxonomy, and linguistics.
Constraints System
Content constraints
control the scope and depth of output content by specifying certain conditions (Zhang et al. 2023), and can be divided into lexical constraints, element constraints, and semantic constraints based on their granularity. Numerical constraints, which ensure that output content meets length and quantity requirements (Yao et al. 2023), can be classified into word-level, sentence-level, paragraph-level, document-level based on the objects involved in the planning. Stylistic constraints impart a unique flavor and color to the output, revealing the author’s traits and chosen social objectives (Tsai et al. 2021), can be subdivided into tonal, formal, audience, and authorial style constraints based on the perspective of application. Format constraints (Tang et al. 2023) standardize expression to guide the generation of complex content and can be categorized into fundamental, bespoke, and specialized scenario constraints based on their usage scenarios. Linguistic constraints (Zhou et al. 2023b) adapt to various scenarios by controlling internal features and logic, grouped into Pragmatic, Syntactic, Morphological, Phonological, and other constraints. Situation constraints (Liu et al. 2023) guide response appropriateness through background or situational parameters, can be classified into role-based, task-specific, and complex contextual constraints. Example constraints regulate new responses by leveraging the intrinsic patterns established by a limited set of samples, with an emphasis on assessing the model’s proficiency in contextual constraint learning. Inverse constraints narrow the response space through the mechanism of indirect exclusion. Contradictory constraints denote conditions that are mutually exclusive, rendering it impossible for the response to fulfill all requirements concurrently, which are prevalent in online logs and are often easily overlooked. Rule constraints define logic flows or actions and meticulously crafted to standardize the road of responses. Details are in Appendix Tab. 8.
Dataset Construction
To guarantee data quality in terms of authority and thorough coverage, we utilize a collaborative iterative methodology that synergizes expertise with the capabilities of LLMs.
Data Source and Selection
Real-world scenarios and NLP tasks form the foundation for CFBench’s initial instructions. By harnessing advanced LLMs, we rigorously assess each instruction for constraint types and quantities within a predefined system, filtering out those with unreasonable or ineffective constraints. Subsequently, we balance the scenarios and constraint types, resulting in a refined set of 2,000 instructions covering all scenarios and NLP tasks. Prompts for instruction and checklist generation are in Appendix.
Iterative Refinement
Professional annotators carefully review and refine the data, ensuring the rationality of constraints and gold answer. If modifications are needed, instructions are revised, and LLMs generate responses with refined evaluation criteria, repeating this process until satisfactory results are achieved. Ultimately, comprehensive support is formulated for each sample, detailing high-quality instructions, the ideal answer, specific assessment criteria, constraint types, and priority levels, totaling 1000 entries.
Figure 3: The distribution of NLP tasks and domains
Models Easy Set Hard Set Full Set
CSR ISR PSR CSR ISR PSR CSR ISR PSR
GPT-4o 0.956 0.868 0.888 0.816 0.438 0.582 0.886 0.653 0.735
GPT-4-Turbo-20240409 0.924 0.792 0.826 0.783 0.370 0.518 0.853 0.581 0.672
GPT-4-0125-Preview 0.923 0.790 0.826 0.763 0.310 0.468 0.843 0.550 0.647
GPT-3.5-Turbo-1106 0.797 0.520 0.602 0.631 0.176 0.326 0.714 0.348 0.464
Claude-3.5-Sonnet 0.943 0.844 0.882 0.799 0.408 0.564 0.871 0.626 0.723
GLM-4-0520 0.939 0.820 0.852 0.785 0.372 0.536 0.862 0.596 0.694
DeepSeek-V2-0628 0.946 0.830 0.868 0.786 0.350 0.524 0.866 0.590 0.696
ERNIE-4-Turbo-0628 0.930 0.790 0.848 0.772 0.332 0.532 0.851 0.561 0.690
Yi-Large 0.900 0.730 0.786 0.744 0.292 0.460 0.822 0.511 0.623
abab6.5-chat 0.894 0.696 0.766 0.736 0.260 0.452 0.815 0.478 0.609
MoonShot-V1-8k 0.919 0.764 0.812 0.758 0.308 0.464 0.838 0.536 0.638
Llama-3-8B-Instruct 0.656 0.300 0.356 0.562 0.122 0.238 0.609 0.211 0.297
Llama-3-70B-Instruct 0.750 0.422 0.498 0.642 0.178 0.330 0.696 0.300 0.414
DeepSeek-V2-Lite-Chat 0.733 0.382 0.448 0.597 0.148 0.262 0.665 0.265 0.355
YI-1.5-34B-Chat 0.881 0.672 0.740 0.745 0.302 0.474 0.813 0.487 0.607
Qwen1.5-110B-Chat 0.905 0.724 0.792 0.730 0.276 0.438 0.818 0.500 0.615
Qwen2-72B-Instruct 0.944 0.836 0.880 0.791 0.342 0.530 0.867 0.589 0.705
Table 2: The evaluation results of LLMs on CFBench and its splits. Notably, stands for the model supporting mainstream languages excluding Chinese, and represents calling through the API. The bold, underlined, and tilde denote the first, second, and third rankings, respectively.
Dataset Statistics
Overall Statistics
Tab. 1 provides a statistical overview of CFBench, highlighting the substantial differences between the two sets. The Hard Set features more detailed instructions, a greater variety of constraints, and a higher number of constraints, indicating a higher level of complexity compared to the Easy Set. Furthermore, the table demonstrates the diversity and balanced distribution of primary constraint types within CFBench, an area where we significantly outperform other benchmarks. See the Appendix for division details.
Task and Domain Distribution
CFBench covers 20 major real-life domains and includes over 200 common scenarios and 50+ NLP tasks. Fig. 3(a) illustrates the classification of NLP tasks, including four major types: classification, generation, sequence labeling, and sentence relation, along with their corresponding specific tasks. Fig. 3(b) shows the real-life scenario-specific domain distribution, where Employment is the most prevalent category, and the other domains are relatively balanced. Our objective is to balance the real distribution with an equitable distribution. Overall, Fig. 3 illustrates that CFBench has evolved into a comprehensive and well-balanced benchmark.
Comparison with Other Benchmarks
As shown in Tab. 3, we thoroughly compare our benchmark with various relevant ones. In terms of size, our benchmark contains approximately twice the number of samples as others. Notably, FollowBench (Jiang et al. 2023), despite having 820 samples, originally had only one-fifth of this number, with each sample expanded by adding five constraints. From the constraint perspective, CFBench is the only benchmark with systematized constraints, featuring the highest number of primary constraint types and ensuring a balanced consideration of these types. Regarding evaluation, InFoBench (Qin et al. 2024) also sets response criteria for each sample, while CFBench distinguishes itself by incorporating user demand prioritization and emphasizing objectivity.
Evaluation Protocol
Evaluation Criteria
We breaking down instructions into multiple simple, independent checkpoints to ensure evaluation accuracy, inspiration was drawn from DRFR (Qin et al. 2024). This approach addresses the challenge of evaluating entire responses, especially for complex instructions with multiple constraints. Unlike DRFR, our method emphasizes defining ideal response characteristics and critical evaluation points. The previous sections detailed the checklist generation process, a key part of our evaluation criteria. Furthermore, we employ the state-of-the-art LLM, GPT-4o, as the evaluation model. By repeatedly feeding it the instruction, test model response, and checklist with a carefully tuned prompt, we ensure that the judged response fully meets the judgement format check. This iterative process aims to maximize confidence in our evaluation. The specific evaluation prompt is in the Appendix.
Evaluation Metrics
Aligned with different perspectives, we define the Constraint Satisfaction Rate (CSR), Instruction Satisfaction Rate (ISR) as follows:
$\mathrm{CSR}=\frac{1}{m}\sum_{\_}{i=1}^{m}(\frac{1}{n_{\_}i}\sum_{\_}{j=1}^{n_{\_}i}s_{\_}i^{j})$ (1)
$\mathrm{ISR}=\frac{1}{m}\sum_{\_}{i=1}^{m}s_{\_}i$ (2)
where $s_{\_}i^{j}$ = 1 if the $j$-th constraint of $i$-th instruction is satisfied and $s_{\_}i^{j}$ = 0 otherwise. $s_{\_}i$ = 1 indicates that all constraints in the $i$-th instruction are satisfied and $s_{\_}i$ = 0 otherwise. The requirements Priority Satisfaction Rate (PSR) is defined as follows:
$\mathrm{PSR}=\frac{1}{m}\sum_{\_}{i=1}^{m}(PSR_{\_}i)$ (3)
Let the average score for secondary requirements be $A$. When all primary requirements are met, $score=0.5+0.5\times A$. If $score>0.8$, then $PSR_{\_}i=1$; otherwise, $PSR_{\_}i=0$, especially when any primary requirement is not met. The threshold of $0.8$ is based on user feedback, reflecting tolerance for LLMs adhering to constraints. Overall, CSR, ISR, and PSR reflect different levels of user perception from multiple perspectives, including constraints, instructions, and requirement priorities.
Benchmarks Data Quality Evaluation
Num. Type. Syst. Prio. Meth.
IFEval 541  4 [Uncaptioned image]
CELLO 523 4 [Uncaptioned image]
FollowBench 820 5 [Uncaptioned image][Uncaptioned image]
InFoBench 500 5 [Uncaptioned image][Uncaptioned image]
CfBench 1000 10 [Uncaptioned image][Uncaptioned image]
Table 3: Detailed Comparison of Relevant Benchmarks. represents our constraint system. ’Num.’, ’Type.’, ’Syst.’, ’Prio.’, and ’Meth.’ denote the number of samples, primary constraint types, presence of a constraint system, requirement prioritization, and evaluation method.
Data Quality
Quality Evolution
To enhance the quality of CFBench, we invested considerable effort and financial resources. First, in the instruction generation phase, we utilized multiple advanced LLMs, such as GPT-4 and Claude, to generate diverse instructions and responses for annotator candidates. Second, we implemented a stringent manual annotation process, including annotator training, cross-validation, batch validation, expert team involvement, and iterative refinement of instruction constraints and response quality. We also ensured the objectivity, evaluability, and prioritization of checkpoints. Additionally, we balanced the data for constraint types, scenarios, and NLP task distribution. Detailed information can be found in the Appendix.
Quality Evaluation
[…]
Figure 4: Different mainstream models’ results under primary and secondary constraint categories.
Figure 5: Different mainstream models’ PSR results in real-world domains and NLP task types.
Experiment
Evaluation Settings
We evaluated a total of 50 models that demonstrated exceptional performance on previous benchmarks  (Hendrycks et al. 2020; Cobbe et al. 2021), considering factors such as model size, support for Chinese among mainstream languages, whether they were accessed via API or weights, and whether they were fine-tuned with instruction tuning data. During inference, we set the maximum generation length to 2048, with other parameters using their default values. For evaluation, we used GPT-4o as the judge model, setting the temperature to 0 to ensure deterministic outputs. All model names mentioned correspond to the versions listed in Tab. 7 of the Appendix.
Model MMLU GSM8K CFBench
GPT-4o 88.7 90.5 0.698
Claude-3.5-Sonnet 88.7 96.4 0.691
Qwen2-72B-Instruct 82.3 91.1 0.672
GLM-4 81.5 87.6 0.665
DeepSeek-V2 78.5 79.2 0.665
Qwen1.5-110B-Chat 80.4 - 0.584
Qwen-1.5-72B-Chat 77.5 79.5 0.577
Table 4: Performance Comparison on Benchmarks
Figure 6: Factors Influencing Constraints-Following Performance
Overall Result
Tab. 2 presents the detailed evaluation results of CFBench and its splits for the leading models, using three key metrics. Overall, GPT-4o leads across all splits and metrics, with Claude-3.5-Sonnet close behind. Qwen2-72B-Instruct, GLM4, DeepSeek-V2, and ERNIE4 also performed well, while Yi-large, abab6.5, and MoonShot lagged slightly. Considering the different splits, the Hard split shows a significant drop in metrics, highlighting the clear distinction between Hard and Easy splits. The highest PSR only reached 0.582, indicating room for improvement. Regarding the metrics, CSR is more favorable for weaker models, making it easier to identify issues at a fine-grained level. ISR and PSR highlight differences among stronger models, with ISR being the most challenging and PSR considering user priority and overall satisfaction. Notably, API-accessed models significantly outperformed open-source models accessed via weights, except for Qwen2-72B-Instruct, highlighting a substantial gap in constraint follow for open-source models.
Constraints-categorized Performance
To observe the performance across different constraint types, we calculated satisfaction scores for each constraint for the top 8 LLMs, as shown in Fig. 4. For primary constraints, many models struggled with contradictory constraints, revealing their limitations in handling conflicts. By considering requirement priorities, we captured nuances among multiple constraints, enabling detailed evaluation. GPT-4o excelled across various constraints, while other models alternated in leading different types. For secondary constraints, all models performed poorly in lexical, word, and sentence count constraints but did better in document count and audience style constraints. No single model consistently led across most constraint types. In summary, even the most advanced LLMs have significant room for improvement in certain constraint types, with each model showing specific weaknesses. This provides valuable insights for future model iterations.
Domain and Task-categorized Performance
As depicted in Fig. 5, we evaluate performance across 21 domains and 4 major NLP task types, each with 500 examples from the two main sources of CFBench. For domain performance, employment and psychology require significant attention, while technology and recruitment are strengths for most models. For NLP tasks, GPT-4o excels in sentence relationship tasks, while Qwen2-72B-Instruct is strong in sequence labeling, likely due to its optimization for Chinese. In general, models exhibit different rankings across domains and tasks, indicating no clear absolute leader. Comprehensive improvements are needed for better constraint follow across multiple domains and tasks.
Discussion
Comparisons between Capabilities
Tab. 4 presents a comprehensive comparison of CFBench’s PSR with two of the prominent LLM evaluation benchmarks: MMLU (Hendrycks et al. 2020) and GSM8K (Cobbe et al. 2021). MMLU focuses on knowledge proficiency, while GSM8K emphasizes mathematical ability. GPT-4o ranks first on CFBench but significantly lags behind, ranking third on GSM8K. Qwen1.5-110B-Chat performs worse than DeepSeek-V2 on CFBench but outperforms it on MMLU. Notably, the rankings of LLMs on CFBench do not entirely correspond with those on the other two benchmarks, indicating that CFBench provides a novel perspective for LLMs evaluation.
Factors influencing constraints-following
Previously, we identified a significant gap in LLM constraints following performance, prompting us to further explore the influencing factors. We analyzed the impact of prompt length, number of constraints, constraint types, and primary requirements on evaluation results across five top-performing models and their average values. As shown in Fig. 6, all four factors are positively correlated with the ISR metric, with the number of constraints having the most significant effect. For PSR, the number of constraints and constraint types do not show a completely positive correlation, while the number of primary requirements has a greater influence. Users are more affected by unmet constraints when there are fewer, but become more tolerant of unmet non-primary constraints when there are many.
How to improve constraint-following ability?
In Appendix Tab. 7, we investigated methods to potentially enhance constraint following. Firstly, Supervised Fine-Tuning (SFT) significantly improves performance, with nearly all models that undergo instruction fine-tuning exhibiting substantial improvements in effectiveness, as demonstrated by the Qwen series. Secondly, model size is also an important factor, as evidenced by Qwen2-72B-Instruct showing a 40% relative PSR improvement over Qwen2-7B-Instruct. Additionally, replicating Conifer’s models  (Sun et al. 2024a) reveals that fine-tuning with complex constraint instructions further enhances performance, and recent work has also been directed towards this approach  (He et al. 2024a). Further exploration is intended to be pursued in future work.
Conclusion
This study comprehensively examines the constraints-following capabilities of LLMs. CFBench, a comprehensive benchmark, was introduced with 1000 manually annotated samples covering more than 200 real-world scenarios and over 50 NLP tasks, encompassing a wide range of systematically defined constraint types. Each sample in CFBench includes detailed evaluation criteria, providing metrics that accurately reflect model performance and real user needs across multiple dimensions. Extensive experiments on CFBench revealed significant limitations and challenges that advanced LLMs face in following constraint instructions. Key factors and potential strategies to improve constraint following were also analyzed. In conclusion, CFBench offers a novel perspective for evaluating LLM capabilities, providing new directions for performance assessment and improvement.