### v2

# CFBench: A Comprehensive Constraints-Following Benchmark for LLMs

Abstract
The adeptness of Large Language Models (LLMs) in comprehending and following natural language instructions is critical for their deployment in sophisticated real-world applications. Existing evaluations mainly focus on fragmented constraints or narrow scenarios, but they overlook the comprehensiveness and authenticity of constraints from the user’s perspective. To bridge this gap, we propose CFBench, a large-scale Chinese Comprehensive Constraints Following Benchmark for LLMs, featuring 1,000 curated samples that cover more than 200 real-life scenarios and over 50 NLP tasks. CFBench meticulously compiles constraints from real-world instructions and constructs an innovative systematic framework for constraint types, which includes 10 primary categories and over 25 subcategories, and ensures each constraint is seamlessly integrated within the instructions. To make certain that the evaluation of LLM outputs aligns with user perceptions, we propose an advanced methodology that integrates multi-dimensional assessment criteria with requirement prioritization, covering various perspectives of constraints, instructions, and requirement fulfillment. Evaluating current leading LLMs on CFBench reveals substantial room for improvement in constraints following, and we further investigate influencing factors and enhancement strategies. The data and code will be made available.
CFBench: A Comprehensive Constraints-Following Benchmark for LLMs
Tao Zhang1thanks:  Equal contribution.  Chenglin Zhu211footnotemark: 1  Yanjun Shen [^1] footnotemark: 1  Wenjing Luo1  Yan Zhang1 Hao Liang2Tao Zhang1Fan Yang1Mingan Lin1Yujing Qiao1 Weipeng Chen1Bin Cui2Wentao Zhang2thanks:  Corresponding author.Zenan Zhou122footnotemark: 2 1Baichuan Inc.  2Peking University {zhangtao2, shenyanjun, zhouzenan}@baichuan-inc.com zhuchenglin@stu.pku.edu.cn, wentao.zhang@pku.edu.cn
1 Introduction
Figure 1: Sample data from CFBench. A checklist, constraint type, requirement priority, and satisfaction constitute our evaluation criteria.
Large Language Models (LLMs) have become the cornerstone of numerous cutting-edge research tasks and are widely utilized in real-world scenarios (Brown et al., 2020; Chowdhery et al., 2023; Achiam et al., 2023; Touvron et al., 2023). In real-world scenarios, human instructions are inherently complex and accompanied by explicit constraints, requiring models to both understand intricate requirements and strictly comply with these constraints Yang et al. (2023); Zhong et al. (2021); Mishra et al. (2022); Wei et al. (2021); Sanh et al. (2022). The proficiency of LLMs in comprehending requirements and adhering to natural language constraints is essential, as it ensures tasks are executed precisely and resolved perfectly according to user instructions.
The prevailing method for evaluating a model’s instruction-following ability involves using quantitative programs, human evaluators, or advanced LLMs to assess response quality across single constraints, complex problems, and finite constraints  Zhou et al. (2023a); Wang et al. (2023); Li et al. (2023); Zheng et al. (2024); Xu et al. (2023).  Laskar et al. (2024) underscores the importance of evaluating data quality, highlighting the necessity for real and extensive data distribution, along with its applicability to real-world scenarios.  Sun et al. (2024b) also stresses that realistic evaluation metrics reflect model capabilities and guide iteration. Constraints-following evaluation faces analogous challenges, particularly within complex real-world scenarios, where data sources and contexts are diverse, and where evaluation is both subjective and arduous. Fig. 1, which addresses the aforementioned challenges, presents a sample from CFBench illustrating the Trump assassination event with different colors representing various constraint types. The instruction include multiple constraints, and the evaluation method uses a checklist to break down complex requirements into independent checkpoints, annotating constraint types and priorities. LLMs are then used to assess each checkpoint. For the English-Chinese comparison example, see Appendix Fig. 11. Consequently, we introduce two more profound challenges in constraints-following assessment.
Q1: How to construct high-quality evaluation data? Many studies focus on evaluating single constraint Chen et al. (2022); Tang et al. (2023), lacking comprehensive analysis across diverse constraints. He et al. (2024b) examines LLM performance on complex real-world instructions but neglect constraint diversity and scenario coverage. Jiang et al. (2023) incrementally incorporate fine-grained constraints to craft multi-level instructions. However, with only 75 instances of mixed type, which risks variability due to limited data, and equating difficulty with constraint quantity oversimplifies the task. Recent work focuses on evaluating constraints combinability Wen et al. (2024). To ensure data quality, we systematically categorize constraints by mining real-world online data and using classification, synthesis, and expert design, covering 10 primary categories and over 25 subcategories. We also cross-match these constraints with various domains and scenarios, ensuring balanced representation and expert-validated reasonableness.
Q2: How to evaluate accurately and meticulously? Evaluating LLMs’ adherence to constraints is challenging and typically involves manual, automated, and programmatic assessments using various metrics. Representative work computes outcomes for verifiable instructions using code  Zhou et al. (2023a); He et al. (2024b).   Jiang et al. (2023) uses scripts and constraint-evolution paths to handle diverse challenging instructions, introducing three metrics tailored to the data’s characteristics. The DRFR method decomposes complex constraints into binary judgments, with GPT evaluating each criterion Qin et al. (2024). Indeed, previous work has ensured the feasibility and objectivity of evaluations through various methods, but they have overlooked assessments from the user’s multiple perspectives. We deconstruct complex instructions from the user’s perspective into multiple sub-needs, categorizing them by priority and constraint type, with LLMs evaluating each checkpoint. Furthermore, a multi-dimensional evaluation criteria is proposed using three metrics from the perspectives of constraints, instructions, and requirements priority.
We introduce CFBench, a comprehensive Chinese benchmark designed to thoroughly evaluate the constraint comprehension and following capabilities of LLMs. CFBench comprises 10 primary categories and over 25 secondary subcategories organized through taxonomic and statistical methodologies. CFBench features 1,000 meticulously curated samples spanning more than 200 real-life scenarios across 20 domains and over 50 NLP tasks, enhancing the breadth and generality of the evaluation data. Additionally, we have seamlessly integrated original instructions and constraint types within each sample, paying particular attention to nuanced combinations, ensuring each constraint is credibly and coherently embedded. Finally, extensive experiments and exploratory discussions provide strong support for evaluation and optimization.
Overall, our contributions are mainly four-fold:
  • To the best of our knowledge, we are the pioneers in systematically defining an instruction constraint framework utilizing both taxonomic and statistical methodologies.
  • We introduce CFBench, a meticulously annotated, large-scale, high-quality Chinese benchmark that encompasses a broad spectrum of real-world scenarios and NLP tasks.
  • We propose a multi-dimensional evaluation framework to comprehensively assess model capabilities while prioritizing user-centric needs.
  • We exhaustively evaluated prominent LLMs, uncovering significant deficiencies in constraints following and exploring performance factors and optimization strategies.
Figure 2: The construction pipeline and evaluation sample of CFBench. Initially, it entails the construction of the constraint system, followed by the assembly of the dataset, and culminating in the proposal of a multi-perspective user view evaluation.
2 Related Work
2.1 Instruction Following
Fine-tuning LLMs with annotated instructional data enhances their ability to follow general language instructions  Weller et al. (2020); Sanh et al. (2022). Studies show that more complex or constrained instructions further improve this ability. For instance, six methods to create intricate instructions from a small set of handwritten seed data are proposed Xu et al. (2023), while  Mukherjee et al. (2023) elevate training data complexity by having GPT-4Achiam et al. (2023) generate reasoning steps for simple instructions. The latest work  Sun et al. (2024a); He et al. (2024a); Dong et al. (2024) suggests that increasing the number and variety of constraints can enhance the complexity of instructions, thereby further improving the model’s ability to follow constraint-based instructions.
2.2 Evaluation of Constraints Following
Constraints such as word count, position, topics, and content have garnered significant attention in the field of Controlled Text Generation  Yao et al. (2023); Zhou et al. (2023b)Zhou et al. (2023a) centers on assessing 25 verifiable instructions. Numerous studies have explored the adherence of LLMs to format constraints, such as complex tabular data  Tang et al. (2023) and customized scenario formats  Xia et al. (2024).  Qin et al. (2024) decomposing a single instruction into multiple constraints. He et al. (2024b) gathered constraints from real-world scenarios and developed a sophisticated benchmark using detailed task descriptions and inputs. Jiang et al. (2023) progressively integrates fine-grained constraints to develop multi-level instructions, thereby enhancing complexity across six distinct types. Concurrent work  Wen et al. (2024), constructs a novel benchmark by synthesizing and refining data from the aforementioned benchmarks, with an emphasis on the combinatorial types of constraints. However, previous studies suffered from fragmented constraints, limited scenarios, and misaligned evaluation methods with user perspectives.
Split Set Basic Info Constraints Count
Num. Len. Prim. Cons. Type. C1 C2 C3 C4 C5 C6 C7 C8 C9 C10
Easy Set 500 413 1.69 3.59 2.83 613 214 180 170 134 92 82 95 90 79
Hard Set 500 605 1.98 4.89 3.58 772 345 233 241 168 122 115 145 137 81
Full Set 1000 509 1.84 4.24 3.20 1385 559 413 411 302 214 197 240 227 160
Table 1: CFBench Statistic. The abbreviations of ’Num.’, ’Len.’, ’Prim.’, ’Cons.’, ’Type.’ denote the sample number, average instruction length, average primary requirements number, average constraint number, average constraint type number, respectively. The designations ’C1’-’C10’ denote the Primary Constraint types of content, numerical, style, format, linguistic, situation, example, inverse, contradictory, and rule constraint, respectively.
3 CFBench
As depicted in Fig. 2, the CFBench construction pipeline includes several key components. First, we collect and systematize constraint expressions from real-world scenarios and various NLP tasks. Using this system, we create high-quality evaluation data by combining instructions from these scenarios with advanced LLMs and manual curation. We then introduce innovative multi-perspective evaluation method. Additionally, we conduct a thorough statistical analysis and validate the quality from various angles to highlight reliability and applicability.
3.1 Constraints System
3.1.1 Constraints Collection
We amass a diverse corpus of instructions from real-world scenarios and various NLP tasks  Xia et al. (2024); Li et al. (2024) to ensure a comprehensive system. Initially, we aggregate several million instructions from online logs and NLP tasks, refining these through length filtering and clustering to distill 30,000 high-quality instructions. Utilizing advanced LLM techniques, we extract and expand atomic constraints through evolutionary methods. Using LLMs, we carefully select meaningful atomic constraints, resulting in over 5000 unique constraints.   Domain experts first filter out unreasonable or meaningless atomic constraints and then synthesize these into a structured framework with 10 primary categories and 25 subcategories, guided by principles of statistics, taxonomy, and linguistics.
3.1.2 Constraints System
Content constraints control the scope and depth of output content by specifying certain conditions Zhang et al. (2023), and can be divided into lexical constraints, element constraints, and semantic constraints based on their granularity. Numerical constraints, which ensure that output content meets length and quantity requirements Yao et al. (2023), can be classified into word-level, sentence-level, paragraph-level, document-level based on the objects involved in the planning. Stylistic constraints impart a unique flavor and color to the output, revealing the author’s traits and chosen social objectives Tsai et al. (2021), can be subdivided into tonal, formal, audience, and authorial style constraints based on the perspective of application. Format constraints Tang et al. (2023) standardize expression to guide the generation of complex content and can be categorized into fundamental, bespoke, and specialized scenario constraints based on their usage scenarios.   Linguistic constraints Zhou et al. (2023b) adapt to various scenarios by controlling internal features and logic, grouped into Pragmatic, Syntactic, Morphological, Phonological, and other constraints. Situation constraints Liu et al. (2023) guide response appropriateness through background or situational parameters, can be classified into role-based, task-specific, and complex contextual constraints. Example constraints regulate new responses by leveraging the intrinsic patterns established by a limited set of samples, with an emphasis on assessing the model’s proficiency in contextual constraint learning. Inverse constraints narrow the response space through the mechanism of indirect exclusion.   Contradictory constraints denote conditions that are mutually exclusive, rendering it impossible for the response to fulfill all requirements concurrently, which are prevalent in online logs and are often easily overlooked. Rule constraints define logic flows or actions and meticulously crafted to standardize the road of responses. Details are in Appendix Tab. 8.
3.2 Dataset Construction
To guarantee data quality in terms of authority and thorough coverage, we utilize a collaborative iterative methodology that synergizes expertise with the capabilities of LLMs.
3.2.1 Data Source and Selection
Real-world scenarios and NLP tasks form the foundation for CFBench’s initial instructions. By harnessing advanced LLMs, we rigorously assess each instruction for constraint types and quantities within a predefined system, filtering out those with unreasonable or ineffective constraints. Subsequently, we balance the scenarios and constraint types, resulting in a refined set of 2,000 instructions covering all scenarios and NLP tasks. Prompts and checklist generation are in Appendix.
3.2.2 Iterative Refinement
Professional annotators carefully review and refine the data, ensuring the rationality of constraints and gold answer. If modifications are needed, instructions are revised, and LLMs generate responses with refined evaluation criteria, repeating this process until satisfactory results are achieved. Ultimately, comprehensive support is formulated for each sample, detailing high-quality instructions, the ideal answer, specific assessment criteria, constraint types, and priority levels.
Figure 3: The distribution of NLP tasks and domains
Models Easy Set Hard Set Full Set
CSR ISR PSR CSR ISR PSR CSR ISR PSR
o1-preview 0.926 0.806 0.844 0.814 0.462 0.592 0.870 0.634 0.718
DeepSeek-V3 0.948 0.836 0.864 0.831 0.460 0.616 0.890 0.648 0.740
DeepSeek-R1 0.960 0.874 0.894 0.856 0.524 0.672 0.908 0.699 0.783
GPT-4o 0.956 0.868 0.888 0.816 0.438 0.582 0.886 0.653 0.735
GPT-4-Turbo-20240409 0.924 0.792 0.826 0.783 0.370 0.518 0.853 0.581 0.672
GPT-4-0125-Preview 0.923 0.790 0.826 0.763 0.310 0.468 0.843 0.550 0.647
Claude-3.5-Sonnet 0.943 0.844 0.882 0.799 0.408 0.564 0.871 0.626 0.723
GLM-4-0520 0.939 0.820 0.852 0.785 0.372 0.536 0.862 0.596 0.694
ERNIE-4-Turbo-0628 0.930 0.790 0.848 0.772 0.332 0.532 0.851 0.561 0.690
Yi-Large 0.900 0.730 0.786 0.744 0.292 0.460 0.822 0.511 0.623
MoonShot-V1-8k 0.919 0.764 0.812 0.758 0.308 0.464 0.838 0.536 0.638
Llama-3-8B-Instruct 0.656 0.300 0.356 0.562 0.122 0.238 0.609 0.211 0.297
Llama-3-70B-Instruct 0.750 0.422 0.498 0.642 0.178 0.330 0.696 0.300 0.414
DeepSeek-V2-Lite-Chat 0.733 0.382 0.448 0.597 0.148 0.262 0.665 0.265 0.355
Qwen1.5-110B-Chat 0.905 0.724 0.792 0.730 0.276 0.438 0.818 0.500 0.615
Qwen2-72B-Instruct 0.944 0.836 0.880 0.791 0.342 0.530 0.867 0.589 0.705
Table 2: The evaluation results of LLMs on CFBench and its splits. Notably, stands for the model supporting mainstream languages excluding Chinese, and represents calling through the API. The bold, underlined, and tilde denote the first, second, and third rankings, respectively.
3.3 Dataset Statistics
3.3.1 Overall Statistics
Table 1 provides a statistical overview of CFBench, highlighting substantial differences between the two sets. The Hard Set has more detailed instructions, a greater variety, and a higher number of constraints, indicating higher complexity compared to the Easy Set. The table also shows the diversity and balanced distribution of primary constraint types within CFBench, outperforming other benchmarks. See the Appendix for division details.
3.3.2 Tasks and Domains Distribution
CFBench covers 20 major real-life domains and includes over 200 common scenarios and 50+ NLP tasks. Fig. 3(a) illustrates the classification of NLP tasks, including four major types: classification, generation, sequence labeling, and sentence relation, along with their corresponding specific tasks. Fig. 3(b) shows the real-life scenario-specific domain distribution, where Employment is the most prevalent category, and the other domains are relatively balanced. Our objective is to balance the real distribution with an equitable distribution. Overall, Fig. 3 illustrates that CFBench has evolved into a comprehensive and well-balanced benchmark.
3.3.3 Comparison with Other Benchmarks
As shown in Tab. 3, we thoroughly compare our benchmark with various relevant ones. In terms of size, our benchmark contains approximately twice the number of samples as others. FollowBench Jiang et al. (2023) increases difficulty by adding the number of constraints, but focuses on the incremental increase of a single constraint type. ComplexBench Wen et al. (2024) places more emphasis on the combination relationships between different constraint types, but only designs four types. IFEval Zhou et al. (2023a) focuses on constraints that can be verified, but lacks generalization. Compared to others, CFBench provides comprehensive scenario coverage, diverse systematic constraints, numerous high-quality samples, and multidimensional evaluation. For details, see Appendix Tab. 10, including Case and Features.
3.4 Evaluation Protocol
3.4.1 Evaluation Criteria
We breaking down instructions into multiple simple, independent checkpoints to ensure evaluation accuracy, inspiration was drawn from DRFR Qin et al. (2024). Unlike DRFR, our method emphasizes defining ideal response characteristics and critical evaluation points. The previous sections detailed the checklist generation process, a key part of our evaluation criteria. Furthermore, we employ GPT-4o, as the evaluation model. By repeatedly feeding it the instruction, test model response, and checklist with a carefully tuned prompt, we ensure that the judged response fully meets the judgement format check. This iterative process aims to maximize confidence in our evaluation. The specific evaluation prompt is in the Appendix.
3.4.2 Evaluation Metrics
Aligned with different perspectives, we define the Constraint Satisfaction Rate (CSR), Instruction Satisfaction Rate (ISR) as follows:
$\mathrm{CSR}=\frac{1}{m}\sum_{\_}{i=1}^{m}(\frac{1}{n_{\_}i}\sum_{\_}{j=1}^{n_{\_}i}s_{\_}i^{j})$ (1)
$\mathrm{ISR}=\frac{1}{m}\sum_{\_}{i=1}^{m}s_{\_}i$ (2)
where $s_{\_}i^{j}$ = 1 if the $j$-th constraint of $i$-th instruction is satisfied and $s_{\_}i^{j}$ = 0 otherwise. $s_{\_}i$ = 1 indicates that all constraints in the $i$-th instruction are satisfied and $s_{\_}i$ = 0 otherwise. The requirements Priority Satisfaction Rate (PSR) is defined as follows:
$\mathrm{PSR}=\frac{1}{m}\sum_{\_}{i=1}^{m}(PSR_{\_}i)$ (3)
Let the average score for secondary requirements be $A$. When all primary requirements are met, $score=0.5+0.5\times A$. If $score>0.8$, then $PSR_{\_}i=1$; otherwise, $PSR_{\_}i=0$, especially when any primary requirement is not met. The threshold of $0.8$ is based on user feedback, reflecting tolerance for LLMs adhering to constraints. Overall, CSR, ISR, and PSR reflect different levels of user perception from multiple perspectives, including constraints, instructions, and requirement priorities.
Benchmarks Data Quality Evaluation
Num. Type. Syst. Prio. Meth.
IFEval 541  4 [Uncaptioned image]
CELLO 523 4 [Uncaptioned image]
FollowBench 820 5 [Uncaptioned image][Uncaptioned image]
InFoBench 500 5 [Uncaptioned image][Uncaptioned image]
FoFoBench 494 1 [Uncaptioned image]
ComplexBench 1150 4 [Uncaptioned image]
CFBench 1000 10-25 [Uncaptioned image][Uncaptioned image]
Table 3: Detailed Comparison of Relevant Benchmarks. represents our constraint system. ’Num.’, ’Type.’, ’Syst.’, ’Prio.’, and ’Meth.’ denote the number of samples, primary constraint types, presence of a constraint system, requirement prioritization, and evaluation method.
3.5 Data Quality
3.5.1 Quality Evolution
To enhance the quality of CFBench, we invested considerable effort and financial resources. First, in the instruction generation phase, we utilized multiple advanced LLMs, such as GPT-4 and Claude, to generate diverse instructions and responses for annotator candidates. Second, we implemented a stringent manual annotation process, including annotator training, cross-validation, batch validation, expert team involvement, and iterative refinement of instruction constraints and response quality. We also ensured the objectivity, evaluability, and prioritization of checkpoints. Additionally, we balanced the data for constraint types, scenarios, and NLP task distribution. Detailed information can be found in the Appendix.
3.5.2 Quality Evaluation
To investigate CFBench’s quality, we randomly selected 100 samples for assessment. Three professional data inspectors evaluated them, resulting in high-quality rates of 94% for instructions, 94% for gold answers, and 93% for checklists (see Appendix Table 5). Additionally, three experts rated Qwen2-7B-Instruct outputs on a 0-1 scale. The kappa coefficient between GPT-4o PSR and expert evaluations was 0.77, highlighting the effectiveness of the PSR evaluation method and metrics, even for smaller models. Further details are in Appendix Table 6.
Figure 4: Different mainstream models’ results under primary and secondary constraint categories.
Figure 5: Different mainstream models’ PSR results in real-world domains and NLP task types.
4 Experiment
4.1 Evaluation Settings
We evaluated 50+ top-performing models from previous benchmarks Hendrycks et al. (2020); Cobbe et al. (2021), considering factors like model size, Chinese language support, access via API or weights, and fine-tuning with instruction data. During inference, we set the maximum generation length to 2048 and used default values for other parameters. For evaluation, we used GPT-4o as the judge model with a temperature of 0 for deterministic outputs.
Figure 6: Factors Influencing Constraints-Following Performance
4.2 Overall Result
Tab. 2 presents CFBench evaluation results for leading models. DeepSeek-R1 DeepSeek-AI et al. (2024) leads overall, followed by GPT-4o and DeepSeek-V3 DeepSeek-AI et al. (2025) in third. Claude-3.5-Sonnet Anthropic (2024) and Qwen2-72B-Instruct Yang et al. (2024) performed well, though DeepSeek-V3 and Qwen2-72B-Instruct showed slight drops in the Full Set. The highest PSR in the Hard Set was 0.582, indicating room for improvement.
While CSR favors weaker models, ISR and PSR highlight differences in stronger models. API-accessed models like GPT-4o outperformed most open-source models, though DeepSeek-V3 and Qwen2-72B-Instruct performed well among open-source models.
Model MMLU GSM8K CFBench
GPT-4o 88.7 90.5 0.698
Claude-3.5-Sonnet 88.7 96.4 0.691
Qwen2-72B-Instruct 82.3 91.1 0.672
DeepSeek-V2 78.5 79.2 0.665
Qwen1.5-110B-Chat 80.4 - 0.584
Qwen1.5-72B-Chat 77.5 79.5 0.577
Table 4: Performance Comparison on Benchmarks
4.3 Constraints-categorized Performance
To assess performance across different constraint types, we calculated satisfaction scores for the top 8 LLMs (see Fig. 4). Many models struggled with contradictory constraints, highlighting their limitations. GPT-4o excelled across various constraints, while other models alternated in leading different types. For secondary constraints, all models performed poorly in lexical, word, and sentence count constraints but did better in document count and audience style constraints. No single model consistently led across most constraint types. In summary, even the most advanced LLMs have significant room for improvement, with each model showing specific weaknesses, providing valuable insights for future iterations.
4.4 Domain and Task-categorized Performance
As depicted in Fig. 5, we evaluate performance across 21 domains and 4 major NLP task types, each with 500 examples from the two main sources of CFBench. For domain performance, employment and psychology require significant attention, while technology and recruitment are strengths for most models. For NLP tasks, GPT-4o excels in sentence relationship tasks, while Qwen2-72B-Instruct is strong in sequence labeling, likely due to its optimization for Chinese. In general, models exhibit different rankings across domains and tasks, indicating no clear absolute leader. Comprehensive improvements are needed for better constraint follow across multiple domains and tasks.
5 Discussions
5.1 Comparisons between Capabilities
Table 4 presents a comprehensive comparison of CFBench’s PSR with two prominent LLM evaluation benchmarks: MMLU Hendrycks et al. (2020) and GSM8K Cobbe et al. (2021). MMLU focuses on knowledge proficiency, while GSM8K emphasizes mathematical ability. GPT-4o ranks first on CFBench but significantly lags behind, ranking third on GSM8K. Qwen2-72B-Instruct performs worse than DeepSeek-V2 on CFBench but outperforms it on MMLU. Notably, the rankings of LLMs on CFBench do not entirely correspond with those on the other two benchmarks, indicating that CFBench provides a novel perspective for LLMs evaluation.
5.2 Factors influencing constraints-following
Previously, we identified a significant gap in LLM constraints following performance, prompting us to further explore the influencing factors. We analyzed the impact of prompt length, number of constraints, constraint types, and primary requirements on evaluation results across five top-performing models and their average values. As shown in Fig. 6, all four factors are positively correlated with the ISR metric, with the number of constraints having the most significant effect. For PSR, the number of constraints and constraint types do not show a completely positive correlation, while the number of primary requirements has a greater influence. Users are more affected by unmet constraints when there are fewer, but become more tolerant of unmet non-primary constraints when there are many.
5.3 How to improve constraint-following ability?
In Appendix Tab. 7, we investigated methods to potentially enhance constraint following. Firstly, Supervised Fine-Tuning (SFT) significantly improves performance, with nearly all models that undergo instruction fine-tuning exhibiting substantial improvements in effectiveness, as demonstrated by the Qwen series. Secondly, model size is also an important factor, as evidenced by Qwen2-72B-Instruct showing a 40% relative PSR improvement over Qwen2-7B-Instruct. Additionally, replicating Conifer’s models  Sun et al. (2024a) reveals that fine-tuning with complex constraint instructions further enhances performance, and recent work has also been directed towards this approach  He et al. (2024a). Further exploration is intended to be pursued in future work.
6 Conclusion
This study comprehensively examines the constraints-following capabilities of LLMs. CFBench, a comprehensive benchmark, was introduced with 1000 manually annotated samples covering more than 200 real-world scenarios and over 50 NLP tasks, encompassing a wide range of systematically defined constraint types. Each sample in CFBench includes detailed evaluation criteria, providing metrics that accurately reflect model performance and real user needs across multiple dimensions. Extensive experiments on CFBench revealed significant limitations and challenges that advanced LLMs face in following constraint instructions. Key factors and potential strategies to improve constraint following were also analyzed, and numerous insightful findings can provide valuable guidance for the optimization of LLMs’ performance. In conclusion, CFBench offers a novel perspective for evaluating LLM capabilities, providing new directions for performance assessment and improvement.
7 Limitations
7.1 Experimental Setup
This study primarily focuses on models with strong Chinese language capabilities, lacking a comprehensive survey of a broader range of English models. Additionally, while we conducted preliminary analyses on the differences in instruction-following abilities between Chinese and English, a more in-depth comparative study is absent.
7.2 Limited Exploration of Reasoning Models
Currently, deep reasoning models like R1 continue to achieve commendable results. However, there is a lack of in-depth research into these models, particularly concerning the factors that enhance their instruction-following abilities.
7.3 Evaluation Model Bias
The evaluation of models predominantly relies on GPT-4o as the judge model. Future research could explore the impact of different evaluation models on assessing the performance of other models.
8 Ethics Statement
This research adheres to the ethical guidelines set forth by the Association for Computational Linguistics (ACL). We have ensured that all data collection and experimental designs comply with privacy protection and informed consent principles, fully respecting and safeguarding the rights of all participants. Furthermore, we have evaluated the potential societal impacts of our research findings, ensuring that their application does not result in adverse effects on society.
Appendix A Appendix
A.1 Constraint System Construction
The construction of the constraint system commenced with the aggregation of data from diverse real-world scenarios and NLP tasks. This encompassed 800,000 query logs from LLM websites over the preceding six months, alongside over 300,000 data points from various NLP tasks. Instructions that were excessively long or short were filtered out, and a vector clustering deduplication algorithm was employed. This meticulous process culminated in a refined dataset comprising approximately 30,000 instructions. Subsequently, GPT was utilized to extract constraint atoms from these instructions, thereby ensuring the comprehensiveness of the constraint system. The prompt employed for GPT-4 extraction, as illustrated in Fig. 7, resulted in the identification of approximately 5,000 unique atomic constraint instructions. Three seasoned experts meticulously refined these into 1,000 significant atomic constraints. By integrating statistical analysis, classification, and linguistic principles, a hierarchical constraint system was developed using Top-Down Organization and Bottom-Up Synthesis methodologies. This system comprises 10 primary categories and 25 secondary categories. The system comprehensively categorizes all types of constraints, ensuring that nearly all specific constraints can be systematically classified within its framework. Detailed information regarding the constraint system is presented in Tab. 8.
A.2 Dataset Construction
We adopted an innovative Human-Machine Collaborative Iterative Construction approach to ensure the highest quality of data. This method involved leveraging advanced LLMs to augment original instructions with additional constraints and generate corresponding responses. These responses were meticulously reviewed for constraint validity, followed by the creation of detailed checklists for each example. Multiple experts participated in this iterative process, continuously refining the outputs by addressing issues encountered by the LLMs and regenerating or manually correcting any substandard samples. The prompts used for GPT to enhance constraints and generate checklists are illustrated in Fig. 8 and Fig. 10. Due to the limited attention given to real-life scenarios, we have meticulously organized and covered 20 domains and over 200 scenarios in our CFBench system, as detailed in Tab. 9. In the end, we gathered 1,000 high-quality data points: 500 from real-world scenarios and 500 from different NLP tasks. Specifically, we implemented the following steps to enhance data quality for manual annotations.
A.2.1 Annotator Training
We sourced annotation contractors from the public and selected 21 candidates for training by seasoned data scientists. After a one-week training program, the annotators engaged in multiple rounds of trial annotations, which were then assessed by data experts. From these assessments, the 9 annotators demonstrating the highest accuracy were selected for this dataset.
A.2.2 Cross-Validation
To reduce the likelihood of missed and incorrect annotations, we implemented an inter-annotator validation process. Three annotators independently reviewed the labeled instructions, responses, and evaluation criteria, achieving a notable agreement rate of 94%. Any discrepancies that emerged were resolved through expert adjudication, ensuring both consistency and accuracy.
A.2.3 Batch Validation
Due to the substantial size of the dataset, it was systematically divided for processing. Following a phased improvement approach, the initial batch sizes were set at 50, 100, and 200, gradually increasing to 400 for later batches. After the annotation process, 50% of the dataset was randomly selected for contractor review, while 20% of the dataset was examined by experts.
A.2.4 Data Split
We used a voting mechanism involving experts and ten models, including GPT-4o and Claude-3.5-Sonnet, to partition 10,000 CFBench entries into ’easy’ and ’hard’ categories. The ’hard’ category includes entries where multiple models struggle with PSR performance and are also challenging for humans, as verified by experts.
Figure 7: Prompt Template Atomic Constraint Extract
Figure 8: Prompt Template for Adding Atomic constraint
A.3 Evaluation Method and Metric
The state-of-the-art GPT-4o model was employed as the judge to perform binary scoring (0 or 1) for each checkpoint in the checklist. The specific evaluation prompt is illustrated in Fig. 9. The Requirement Priority-Satisfaction Ratio (PSR) was proposed as an evaluation metric that simultaneously considers the prioritization of user requirements and satisfaction levels. PSR is calculated by first ensuring that all primary requirements are met. Subsequently, the satisfaction score is determined by averaging the fulfillment of the remaining constraints to obtain $A$. The final satisfaction score is then calculated using the formula $0.5+0.5*A$. If the final score exceeds $0.8$, PSR is set to $1$. The threshold of $0.8$ was established based on the average satisfaction levels derived from multiple users’ feedback on the responses to the instructions.
A.4 Quality Assessment
We employed multiple methods to validate the quality of the benchmark on a randomly selected set of 100 samples. First, we engaged three experts to independently evaluate the quality of each sample’s instruction, response, and criteria. The average quality rate determined by the three experts was consistently above 90%, as detailed in Tab. 5. To further validate the effectiveness of our proposed evaluation metric, PSR, we had the same three experts score the responses of Qwen2-7B-Instruct on these 100 cases using a 0-1 scale. Simultaneously, we utilized GPT-4o to directly score the responses, referred to as GPT-4o PSR. By calculating the kappa coefficient, we found a strong agreement between our proposed PSR evaluation metric and the human experts’ assessments. The detailed results are presented in Tab. 6. Kappa coefficient scores are interpreted as follows: below 0.2 indicates slight agreement, 0.21 to 0.40 indicates fair agreement, 0.41 to 0.60 indicates moderate agreement, 0.61 to 0.80 indicates substantial agreement, and 0.81 to 1.00 indicates almost perfect agreement.
Figure 9: Prompt Template for Evaluation
Set Instruction Gold Ans CheckList
Easy Set 0.96 0.94 0.93
Hard Set 0.92 0.95 0.93
All Set 0.94 0.94 0.93
Table 5: The High-Quality Rate of 100 selected Samples
Set Easy Set Hard Set Full Set
Avg.Expert 1 1 1
GPT-4o DS 0.58 0.61 0.60
GPT-4o PSR 0.76 0.77 0.77
Qwen2-72B-Inst. PSR 0.70 0.73 0.72
Table 6: The kappa coefficient between expert evaluations and various assessment methods
A.5 Experimental Setup and Results
A.5.1 Experiment Setting
We evaluated the most popular Large Language Models (LLMs), with the majority of these models being developed by companies based in China, primarily to accommodate our CFBench’s focus on the Chinese language. Among the 50 evaluated models, they can be categorized into two groups based on their access method: API-based and open-source weight-based models. It is worth noting that the Llama series models do not primarily support the Chinese language, which results in noticeably lower performance. Both conifer-base and conifer-test are based on the Mistral-7B foundational model. Llama-3-8B-Instruct-CN and Llama-3-70B-Instruct-CN respectively represent Llama-3-8B-Instruct-Chinese and Llama-3-70B-Instruct-Chinese, both of which have undergone Chinese SFT (Supervised Fine-Tuning). For the base models, we used a 3-shot approach to ensure a fair evaluation. The complete list of evaluated models can be found in Tab. 7.
A.5.2 Explanation of Results
GPT-4o and Claude3.5-Sonnet have demonstrated near-absolute leadership, achieving outstanding performance across various metrics and categories. Similarly, models such as GLM-4-0510, ERNIE-4-Bot-0613, ERNIE-4-Turbo-0628, DeepSeek-V2-0628, and Qwen2-72B-Instruct have also exhibited strong capabilities. Many models that support less mainstream Chinese languages performed significantly worse, which is unfair to them and only serves to illustrate their relative rankings. This also confirms that performance are highly correlated with language, especially within the scope of language constraints. From the perspective of open-source versus closed-source models, open-source models have generally achieved comprehensive success. However, Qwen2-72B-Instruct, as an open-source model, also demonstrated notable constraint-following capabilities. Regarding model size, within the Qwen series, performance metrics clearly improve with increasing model size. Additionally, models that have undergone Supervised Fine-Tuning (SFT) show significantly enhanced instruction-following capabilities. The complete evaluation results and rankings can be found in Tab. 7.
Models Easy Set Hard Set Full Set
CSR ISR PSR CSR ISR PSR CSR ISR PSR
GPT-4o 0.956 0.868 0.888 0.816 0.438 0.582 0.886 0.653 0.735
GPT-4-Turbo-20240409 0.924 0.792 0.826 0.783 0.370 0.518 0.853 0.581 0.672
GPT-4-0125-Preview 0.923 0.790 0.826 0.763 0.310 0.468 0.843 0.550 0.647
GPT-3.5-Turbo-1106 0.797 0.520 0.602 0.631 0.176 0.326 0.714 0.348 0.464
Claude-3.5-Sonnet 0.943 0.844 0.882 0.799 0.408 0.564 0.871 0.626 0.723
GLM-4-0520 0.939 0.820 0.852 0.785 0.372 0.536 0.862 0.596 0.694
DeepSeek-V2-0628 0.946 0.830 0.868 0.786 0.350 0.524 0.866 0.590 0.696
ERNIE-4-Turbo-0628 0.930 0.790 0.848 0.772 0.332 0.532 0.851 0.561 0.690
ERNIE-4-Bot-0613 0.929 0.792 0.832 0.779 0.338 0.518 0.854 0.565 0.675
ERNIE-3.5-0613 0.901 0.720 0.772 0.758 0.302 0.482 0.830 0.511 0.627
Yi-Large 0.900 0.730 0.786 0.744 0.292 0.460 0.822 0.511 0.623
abab6.5-chat 0.894 0.696 0.766 0.736 0.260 0.452 0.815 0.478 0.609
MoonShot-V1-8k 0.919 0.764 0.812 0.758 0.308 0.464 0.838 0.536 0.638
Vicuna-7B-V13 0.563 0.206 0.262 0.468 0.100 0.168 0.516 0.153 0.215
Vicuna-33B-V13 0.621 0.270 0.352 0.527 0.110 0.196 0.574 0.190 0.274
Vicuna-13B-V13 0.605 0.248 0.302 0.503 0.100 0.178 0.554 0.174 0.240
Llama-2-7B-Chat 0.5268 0.198 0.250 0.448 0.096 0.152 0.487 0.147 0.201
Llama-2-13B-Chat 0.574 0.242 0.280 0.488 0.094 0.178 0.531 0.168 0.229
Llama-3-8B-Instruct 0.656 0.300 0.356 0.562 0.122 0.238 0.609 0.211 0.297
Llama-3-70B-Instruct 0.750 0.422 0.498 0.642 0.178 0.330 0.696 0.300 0.414
Mistral-7B-Instruct-V03 0.227 0.072 0.086 0.148 0.008 0.022 0.188 0.040 0.054
Conifer-Base 0.510 0.184 0.232 0.300 0.018 0.048 0.405 0.101 0.140
Conifer-Test 0.559 0.215 0.255 0.328 0.102 0.156 0.443 0.159 0.206
BaiChuan-13B-Chat 0.630 0.306 0.366 0.521 0.114 0.196 0.575 0.210 0.281
BaiChuan2-13B-Chat 0.669 0.348 0.418 0.547 0.134 0.226 0.608 0.241 0.322
Llama-3-8B-Instruct-CN 0.743 0.458 0.510 0.627 0.162 0.314 0.685 0.310 0.412
Llama-3-70B-Instruct-CN 0.756 0.482 0.536 0.636 0.190 0.322 0.696 0.336 0.429
DeepSeek-7B-Chat 0.695 0.378 0.442 0.580 0.150 0.270 0.638 0.264 0.356
DeepSeek-V2-Lite-Chat 0.733 0.382 0.448 0.597 0.148 0.262 0.665 0.265 0.355
DeepSeek-67B-Chat 0.802 0.516 0.578 0.662 0.180 0.350 0.732 0.348 0.464
InternLM2-Chat-7B 0.767 0.452 0.538 0.625 0.172 0.320 0.696 0.312 0.429
GLM-4-9B-Chat 0.885 0.678 0.742 0.742 0.288 0.450 0.813 0.483 0.596
YI-1.5-34B-Chat 0.881 0.672 0.740 0.745 0.302 0.474 0.813 0.487 0.607
Qwen1.5-4B 0.454 0.170 0.198 0.376 0.074 0.116 0.415 0.122 0.157
Qwen1.5-4B-Chat 0.652 0.310 0.362 0.536 0.104 0.198 0.594 0.207 0.280
Qwen1.5-7B 0.473 0.176 0.212 0.400 0.090 0.142 0.437 0.133 0.177
Qwen1.5-7B-Chat 0.799 0.534 0.592 0.654 0.194 0.338 0.726 0.364 0.465
Qwen1.5-14B 0.498 0.228 0.280 0.430 0.110 0.176 0.464 0.169 0.228
Qwen1.5-14B-Chat 0.822 0.558 0.626 0.671 0.202 0.370 0.746 0.380 0.498
Qwen1.5-32B 0.647 0.336 0.408 0.528 0.132 0.224 0.587 0.234 0.316
Qwen1.5-32B-Chat 0.883 0.678 0.744 0.704 0.228 0.412 0.793 0.453 0.578
Qwen1.5-72B 0.627 0.324 0.380 0.556 0.148 0.248 0.591 0.236 0.314
Qwen1.5-72B-Chat 0.896 0.710 0.776 0.730 0.254 0.436 0.813 0.482 0.606
Qwen1.5-110B-Chat 0.905 0.724 0.792 0.730 0.276 0.438 0.818 0.500 0.615
Qwen2-0.5B-Instruct 0.446 0.150 0.172 0.393 0.070 0.110 0.419 0.110 0.141
Qwen2-1.5B-Instruct 0.607 0.250 0.316 0.496 0.104 0.168 0.551 0.177 0.242
Qwen2-7B 0.576 0.260 0.316 0.478 0.120 0.192 0.527 0.190 0.254
Qwen2-7B-Instruct 0.835 0.584 0.642 0.682 0.198 0.362 0.758 0.391 0.502
Qwen2-72B 0.711 0.424 0.484 0.568 0.170 0.274 0.640 0.297 0.379
Qwen2-72B-Instruct 0.944 0.836 0.880 0.791 0.342 0.530 0.867 0.589 0.705
Table 7: The complete evaluation results and rankings of CFBench and its respective subsets. Notably, stands for the model supporting mainstream languages excluding Chinese, and represents calling through the API. The bold, underlined, and tilde denote the first, second, and third rankings, respectively. Llama-3-8B-Instruct-CN and Llama-3-70B-Instruct-CN respectively represent Llama-3-8B-Instruct-Chinese and Llama-3-70B-Instruct-Chinese, both of which have undergone Chinese SFT (Supervised Fine-Tuning). Both coniferbase and conifertest are based on the Mistral-7B foundational model. For the base model, we used a 3-shot approach for generation.
Primary Secondary Definition Example
Content Constraint Lexical Mandatory use of specific terms or symbols, including their inclusion and precise placement. …must include the word "beautiful."
Element Mandates for including specific elements or concepts in responses, reflecting a scenario or object. …highlights the Great Wall.
Semantic Directives on thematic content, perspective, or tone, emphasizing response significance. Write a poem about London.
Numerical Constraint Word Count Limit the number of words or tokens. A 50-word poem.
Sentence Count Limit the number of sentences. … three sentences.
Paragraph Count Limit the number of paragraphs. divided into 3 sections.
Document Count Limit the number of documents. … list 3 articles.
Stylistic Constraint Tone and emotion The emotional tone must adhere to standards such as seriousness, anger, joy, humor, and politeness. Write a letter in an angry and sarcastic tone.
Form and style Text expression standards ensure alignment with specific stylistic criteria in both presentation and perception. Write a passage in an encyclopedic style.
Audience-specific Text should be tailored to specific audiences, ensuring clarity and relevance for children, students, or specialized groups. Write a pome for a 6-year-old.
Authorial style Texts should emulate the styles of authors like Shakespeare to achieve artistic effects or depth. Write a passage the style of Shakespeare.
Format Constraint Fundamental Widely accepted and utilized standard formats, including JSON, XML, LaTeX, HTML, Table, and Markdown. Extract keywords and output in JSON format.
Bespoke Protocols for information expression tailored to specific needs,including paragraphing, headings, text emphasis, examples, and bullet points. Summarize the main idea and output in unordered list format.
Specialized Formatting standards tailored for specialized applications or domains. Conform to electronic medical record format.
Linguistic Constraint Pragmatic Contextual language study, encompassing speech acts, implicature, discourse, dialects, sociolects, and language policy. Output in English, in classical Chinese style.
Syntactic Sentence structure, including phrases, constituents, subordinate clauses, ba-constructions, and imperatives. Use imperatives with nouns and verb phrases.
Morphological The internal structure and formation rules of words, including roots, affixes, and morphological changes. Output all content in lowercase English.
Phonological Study on phonological structures:phonemes, allophones, pitch, duration, and intensity. Single-rhyme tongue twisters.
Situation Constraint Role-based Simulating characters based on context, emulating their traits, language, and behaviors. You are Confucius, how do you decide?
Task-specific Offer tailored solutions based on a nuanced understanding of situational demands. Must work from home, how to report?
Complex context Reasoning and problem-solving within intricate and multifaceted contexts. 4 on the left, 10 total, which from right?
Example Constraint - Regulate new responses by leveraging intrinsic patterns from a limited set of samples. Example:input:xxx, output:{…}; input:xx, output?
Inverse Constraint - Narrow the response space through inverse constraints and indirect exclusion. Prohibited from answer political topics.
Contradictory Constraint - Mutually exclusive constraints prevent fulfilling all requirements concurrently. Write a five-character quatrain, 1000 words.
Rule Constraint - Standardize the road of responses through meticulously crafted logic flows or actions. Each answer adds 1, 1+1=3, then 2+3=?
Table 8: Constraint System of CFBench
Domain Scenarios List
Healthcare Symptom Consultation Diagnostic Explanation Medication Guidance Procedures
Wellness Medical Info Guidelines Inquiry Public Health
Medical Education
Education Teaching Methods Resource Access Curriculum Design Communication
Academic Counseling Mental Health Support Tutoring Subject Q&A
Reports Interests
Finance Market Research Stock Analysis Investment Analysis Personal Finance
Corporate Tax Insurance Management Corporate Financing Compliance & Risk
Financial Education Product Development Customer Service Financial Reports
Regulatory Analysis
Legal Legal Education Legal Consultation Document Review Case Analysis
Statute Explanation Regulation Analysis IP Management Legal Training
Case Management Compliance & Risk
Media Content Creation Information Analysis Marketing & Promotion News Reporting
Tourism Travel Consultation Itinerary Planning Route Introduction
Recruitment JD Writing & Analysis Resume Creation Resume Screening Interview Preparation
Interview Evaluation Career Planning Offer Comparison Communication Skills
Performance Review
Gov Affairs Policy Research Public Education Service Guide Public Services
Document Writing Content Review Business Procedures Civil Servant Training
Emergency Management
Real Estate Purchase Planning Market Trends Property Policies Development
Leasing Property Valuation Amenities Financial Services
Property Description Content Creation Sales & Marketing Qualifications
Renovation
Automotive Marketing & Sales Driving & Safety Customer Experience Model Consultation
Model Comparison Loan Calculation Insurance Evaluation Claims Assessment
Car Reviews Maintenance & Repair
Psychology Romance Family Friendship Workplace
Self & Health Social Sexuality & Gender Life Stages
Organizational Client Relations Crisis Intervention Public Psychology
Internet Business Analysis Product Design User Research Coding & Debugging
Product Testing Data Management Cybersecurity Computer Q&A
Internet News Marketing Operations UI/UX Design
Spirituality Beliefs & Rituals Divination Feng Shui Astrology
Metaphysics Spirituality Healing Content Review
Sports Training Goal Setting Nutrition Workout Plans
Equipment & Tech Performance Injury Care Specialized Training
Mental Motivation Data Tracking
Lifestyle Life Tips Shopping Decisions Instant Queries Skincare
Fashion & Styling Naming Recommendations Planning
Socializing Life Creations Q&A
Culturtainment Podcasts & Radio TV & Film Music Literature
Theater & Dance Art Cultural Events Short Videos & Live
Gossip Content Creation
Employment Project Management Translation Office Efficiency Marketing
Administration Customer Service Team Collaboration Collaboration
Cuisine Food & Restaurant Recs Reviews & Feedback Marketing & Promo Food Content
Culinary Culture Recipes & Menus Cooking Techniques Ingredient Prep
Nutrition & Health Food Safety Culinary Training
Gaming Guides Reviews Hardware & Peripherals News
Software & Services Development & Design Operations Mini Games
Search Marketing & Promotion Esports & Tournaments Culture & Education
Technology Reviews Launches Buying Guides Tips & Tricks
Content Creation Product Design Marketing Copy After-sales & Repairs
Table 9: Domain and Scenarios List
Benchmark Case Features CFBench’s Advantages
Complex
Bench
Instruction: Please introduce the following painting.
- If the work contains any animal, the description should be in English.
- Otherwise, the description should be in Chinese.
Response: Painting: "Mona Lisa"
Focus on the combination of instructions, including And/Chain/Selection/Nested.
CFBench supports simplicity in instruction compliance, broad applicability across use cases, and scalability for various tasks.
Follow
Bench
Instruction: Using following words to generate a story.
"needed, money, computer, bought, happy".
Response: In the heart of the city, there was a research facility that desperately needed funding for a crucial project. To gather the required money, ···
Focus on gradual constraint escalation. The more constraints per instruction, the greater the complexity.
CFBench is simple to use, resource-efficient, and scalable for large models.
IFEval Instruction: Given the sentence "Two young boys with toy guns and horns." can you ask a question? Please ensure that your response is in English, and in all lowercase letters. No capital letters are allowed.
Response: what are the two young boys holding?
Verifiable Instructions.
CFBench is easy to implement, general-purpose across models, and scalable for large datasets.
CFBench Instruction: Overview of the Trump Shooting Incident, include the cause, process, aftermath, shooter information and Trump’s feedback. Organize the content using numbered points and bold text. The cause should be in JSON format with event keywords: time, location, and nature. Summarize the process in three sentences. Discuss the impact from both international and domestic perspectives, keeping the total content under 500 words. Finally, summarize this event in a seven-word quatrain.
[…]
Comprehensive Coverage:
Multi-dimensional constraints;
Cross-task evaluation;
Real-world complexity.
Standardized Framework:
Unified metrics;
Cross-model compatibility.
[Uncaptioned image] [Uncaptioned image]
Diverse instruction types,
standardized evaluation,
hierarchical constraints.
Table 10: Benchmark Comparison for Different Instruction Types and Task Complexity
Figure 10: Prompt Template for CheckList Generation
Figure 11: CFBench Example: English-Chinese Comparison, Data Itself in Chinese.